OpenStudy (anonymous):

how many numbers between 8 and 195 are divisible by 6? what is their sum?

OpenStudy (anonymous):

who can help?

OpenStudy (anonymous):

sambhav can u help me pls????

OpenStudy (anonymous):

with the use of arithmetic progression

OpenStudy (anonymous):

pls help me... im confused about this

OpenStudy (anonymous):

the answer is 810

OpenStudy (anonymous):

do u want the explanation ???

OpenStudy (anonymous):

yes sir... can u pls im thanking u in advance

OpenStudy (anonymous):

first digit is 12 last digit divisible by 6 within 195 is 192 so the difference is 6 using formula find out the no. of digit\[n= \frac{ l-a }{ d} +1=\frac{ 192-12 }{ 6 } +1 =31\]

OpenStudy (anonymous):

then using the summation formula sum\[=\frac{ n }{ 2 }\left( 2a+\left( n-1 \right) d\right)\] \[\frac{ 31}{ 2 }(24+(14)6) =3348\] iand apolologies for my wrong answer idid it in a hurry

OpenStudy (anonymous):

is this wrong sir?

OpenStudy (anonymous):

thisone is write i gasve u step by step answer and the firast one which i said'810" was wrong i again apologies for that